# Remarks on Lang’s Conjectural Characterisation of Hyperbolicity of Projective Manifolds

### v1

@@ Acknowledgements
Proceed by contradiction and suppose that $f:\mathbf{C}\longrightarrow X$ is an entire curve. Let $V:=\overline{f(\mathbf{C})}^{\text{Zar}}$ denote its Zariski closure. We choose the entire curve such that the dimension $d\geq 1$ of $V$ is minimal. We may and shall assume that $X$ does not support any rational curves or images from an Abelian variety, and so $d>1$. Let $V^{\text{norm}}$ denote the normalization of $V$. Since $V$ is not uniruled, the Kodaira dimension of $V^{\text{norm}}$ is nonnegative.
The singularities of $V$ can be resolved to produce a smooth model $V^{\text{smth}}$. Running the minimal model program then yields a terminal good minimal model $V^{\text{term}}$ of $V^{\text{smth}}$. The canonical bundle of the terminal model is nef and, hence, semi-ample by the abundance conjecture. Let $\Phi:V^{\text{term}}\longrightarrow W$ denote the Moishezon–Iitaka map. The generic fiber $V_{w}^{\text{term}}=\Phi^{-1}(w)$ have terminal singularities and vanishing Kodaira dimension. Hence, by the abundance conjecture, the canonical bundle of $V_{w}^{\text{term}}$ is torsion. Since $V$ does not contain a Zariski dense set of rational curves, the same is true for $V_{w}^{\text{term}}$. Hence, by Conjecture 1.8, the generic fiber is an étale quotient of an Abelian variety. In particular, $V_{w}^{\text{term}}$ contains Zariski dense entire curves. Since the dimension $d$ is chosen to be minimal, we must have $V_{w}^{\text{term}}\simeq V^{\text{term}}$, and therefore, $W$ is a point, and $V^{\text{term}}$ is an étale quotient of a simple Abelian variety.
We claim that $V^{\text{norm}}\simeq V^{\text{term}}$. Indeed, the fibers of the birational map $\varphi:V^{\text{smth}}\longrightarrow V^{\text{term}}$ are rationally chain connected. Since $V^{\text{norm}}$ does not have any rational curves, it follows that the fibers of $\varphi$ are contained in the fibers of $V^{\text{smth}}\longrightarrow V^{\text{norm}}$. The rational map $\psi:V^{\text{term}}\dashrightarrow V^{\text{norm}}$ is, therefore, regular. Given that $V^{\text{term}}$ is an étale quotient of an Abelian variety $A$, there are no subvarieties of $V^{\text{term}}$ that contract to a point (these would move under the translations of $A$); the map $\psi$ is thus biregular, proving the claim. ∎
@@
In the proof of Theorem 1.1, we only made use of the following weaker form of Conjecture 1.8: Let $V^{\text{norm}}$ be a normal projective variety with no rational curves. Let $V^{\text{term}}$ be a terminal model with torsion canonical bundle, and write $\psi:V^{\text{term}}\dashrightarrow V^{\text{norm}}$ for the rational map. Let $\text{Exc}(\psi)$ and $\Gamma(\psi)$ denote the exceptional locus and graph of $\psi$, respectively. If all rational curves of $V^{\text{term}}$ are contained in the image of $\text{Exc}(\psi)$ under the projection $\Gamma(\psi)\longrightarrow V^{\text{term}}$, then $V^{\text{term}}$ is an étale quotient of an Abelian variety.
In the proof of Theorem 1.1, we showed the following:
Corollary 2.2.
Let $V$ be a normal projective variety with torsion canonical bundle and no rational curves. If Conjecture 1.8 holds, then $V$ is an étale quotient of an Abelian variety.
2.2. Proof of Theorem 1.3.
Proof of Theorem 1.3.
Since $X$ does not contain any rational curves, the canonical bundle $K_{X}$ is nef [34]. Hence, Conjecture 1.6 implies that $K_{X}$ is semi-ample and the smooth fibers of the Moishezon–Iitaka map $\Phi:X\longrightarrow Z$ have torsion canonical bundle. Since the fibers cannot contain rational curves, it follows from Conjecture 1.7 that the smooth fibers are étale quotients of Abelian varieties.
Lemma 2.3.
Let $f:X\longrightarrow Z$ be a fibration from a projective manifold $X$ onto the normal projective complex space $Z$. Assume that the smooth fibers of $f$ are étale quotients of Abelian varieties and that $X$ does not contain any rational curve. Assuming conjecture 1.8, there exists a finite étale cover $u:X^{\prime}\longrightarrow X$ such that if $f^{\prime}:X^{\prime}\longrightarrow Z^{\prime}$, $v:Z^{\prime}\longrightarrow Z$ is the Stein factorization of $f\circ u=v\circ f^{\prime}:X^{\prime}\longrightarrow Z$, then $f^{\prime}:X^{\prime}\longrightarrow X$ is a submersion with Abelian fibers.
Proof.
This is proved in [19] for compact Kähler manifolds with nef cotangent bundle, but without conjecture 1.8. We will show that his proof can be adapted to prove Lemma 2.3. It suffices to show that over the discriminant locus $\text{disc}(f)\subset Z^{\text{reg}}$ of $f$, the fibers are multiple with smooth reduction. Once this has been established, the proof is identical to what is presented in [19]. We note, however, that the proof makes use of [27, Theorem 6.3] deducing the result from the finiteness of the local monodromies around $\text{disc}(f)$ and from the fundamental results of Deligne on the variations of Hodge structures [12].
We thus have to show that if $g:Y\longrightarrow\mathbf{D}$ is a projective fibration from the smooth $Y$ over the unit disk $\mathbf{D}$, with all of its fibers outside of the central one étale quotients of Abelian varieties, if $Y$ does not contain rational curves, and if $K_{Y}$ is torsion, then the central fiber is a multiple of a smooth manifold.
We first recall the following standard assertions:
Lemma 2.4.
Let $V\subset Y$ be an irreducible projective divisor in a complex manifold $Y$, $\nu:V^{\text{norm}}\longrightarrow V$ its normalization and $\delta:V^{\text{smth}}\longrightarrow V^{\text{norm}}$ a desingularization. The canonical bundle $K_{V}:=K_{Y|V}+\mathcal{O}_{Y}(V)$ is defined by adjunction and $K_{V^{\text{norm}}}=j_{*}(K_{(V^{\text{norm}})^{\text{reg}}})$ is the extension of the canonical bundle of its smooth part. Then:
  • (i)
    $K_{V^{\text{norm}}}=\nu^{*}(K_{V})-D$, where $D$ is an effective Weil divisor of $V^{\text{norm}}$ (given by the conductor ideal).
  • (ii)
    $\kappa(V^{\text{smth}})\leq\kappa(V^{\text{norm}})\leq\kappa(V,K_{V})$. [^6] Statement (ii) directly follows from the definitions. Moreover, strict inequalities occur in general: the plane nodal cubic, or the Ueno threefold quotient of $E^{\oplus 3}$ by $\mathbf{Z}_{4}$ acting diagonally on the product, $E$ the Gauss elliptic curve $E$ with complex multiplication of order $4$.
Write $F=\sum_{j=1}^{m}t_{j}F_{j}$ for the decomposition into irreducible components of the scheme-theoretic fiber. By adjunction and up to torsion, we have
$K_{F_{1}}$ $\simeq$ $K_{F}|_{F_{1}}-\left(\sum_{j=2}^{m}\frac{t_{j}}{t_{1}}F_{j}\right)|F_{1}.$
Suppose the reduction $F_{\text{red}}$ is not irreducible, equal to $F_{1}$. Since $K_{Y}|F_{1}$ is torsion, $K_{F_{1}}$ is anti-effective, and Lemma 2.4 implies that $\kappa(F_{1}^{\text{smth}})=-\infty$ for any smooth model $F_{1}^{\text{smth}}$. Since $F$ is not uniruled, however, this cannot be the case, and so $F_{\text{red}}=t_{1}F_{1}$.
It follows that $K_{F_{1}}$ is torsion, and since $F_{1}$ is normal, Lemma 2.4 implies $\kappa(F_{1}^{\text{smth}})=0$. Applying Corollary 2.2, $F_{1}$ is an étale quotient of an Abelian variety, and thus smooth.
We can thus now apply directly [19, Corollary 2.3].
After a finite étale cover of $X$, we may thus assume that $\Phi:X\longrightarrow Z$ is a submersion with all fibers Abelian varieties. We thus get a holomorphic moduli map $\mu:Z\longrightarrow\mathcal{A}_{g,P}$ to the moduli space of Abelian varieties of dimension $g:=\dim(X)-\dim(Z)$ with a suitable polarization. The image is a compact subvariety of general type (this moduli space carries a Hermitian metric with Griffiths semi-negative curvature and negative holomorphic sectional curvature, see [27, Lemma 5.9.3]). Since $Z$ is weakly special, the moduli map $\mu$ is constant, and $\Phi$ is isotrivial; hence, we see that $\kappa(X)=\kappa(Z)$. On the other hand, $\kappa(X)=\dim(Z)$ since $\Phi$ is the Moishezon–Iitaka map of $X$. Therefore, $\kappa(Z)=\dim(Z)$ and $Z$ is of general type and special. Thus $\dim(Z)=0$, and $X$ is the (only) fiber of $\Phi$, hence an Abelian variety (after étale cover). ∎
Remark 2.5.
From Conjecture 1.6 and the preceding arguments, one concludes that $X$ is canonically polarised since $K_{X}$ is big and nef since $X$ does not contain any rational curve and any étale quotient of an Abelian variety (this by Conjecture 1.7). However, a projective manifold of general type without rational curve is canonically polarised by Mori’s theorem [34] and the Kawamata–Shokurov basepoint-free theorem [22].
3. Examples: The Proof of Theorems  1.10 and 1.11.
@@
In this section, we prove Theorems 1.10 and 1.11, producing examples of canonically polarised manifolds with exceptional locus (in the sense of Lang) any given simple Abelian variety, or more generally manifolds with Zariski dense entire curves.
@@
Choose for $V$ a simple Abelian variety. The projective manifold $X$ constructed above is canonically polarised and does not contain any rational curve or elliptic curve if $d\geq 2$. Any entire curve in $X$ is Zariski dense in some of the exceptional divisors $V_{j}\cong V$ of $\alpha:X\longrightarrow X_{0}$ which lie over $\psi^{-1}(v)$ and are the only subvarieties of $X$ not of general type. Moreover, the cotangent bundle of $X$ is nef and big, and the Kobayashi pseudodistance of $X$ is a (non-degenerate) distance on $X\setminus W$, and vanishes on $W$, the union of the $V_{j}^{\prime}s$.
@@
Since $A$ is simple, any of its subvarieties (in particular, $X_{0}$) are of general type. So is thus the manifold $X$. Since neither $A$ nor $V$ carries rational curves, $K_{X}$ is ample (nef by [34], ample by [22]). Since $A$ is simple, $X_{0}$ does not carry any rational curve or Abelian subvariety. Thus $X_{0}$ is hyperbolic, and any entire curve of $X$ must be contained in the exceptional locus of $\alpha:X\longrightarrow X_{0}$, which is the union of the $V_{j}^{\prime}s$. Since each $V_{j}$ is isomorphic to the simple Abelian variety $V$, each entire curve of $X$ is contained and Zariski dense in one of the $V_{j}^{\prime}s$ (by Bloch’s theorem).
We prove that the cotangent bundle of $X$ is nef and big. By [11], Proposition 5, it is sufficient for this to check that $X\subset B=A\times V$ is non-degenerate, that is, that the natural difference map $\delta:X\times X\longrightarrow B$ defined by $\delta(x,y):=x-y$ is generically finite, using the group law on $B$, whatever the origin. Since $A$ and $V$ are simple and not isogeneous (they have different dimensions), the only Abelian subvarieties of $B$ are $A$ and $V$. So we only need to check that for both projections from $B$ to $A$ and $V$, the image of $X$ is either everything or of the same dimension as $X$. But this is obvious since $X$ maps surjectively onto $V$ and birationally to $X_{0}$ in $A$.
We now prove that the Kobayashi pseudodistance on $X$ is a distance on $X\setminus W$. To this end, we observe that $X_{0}\subset A$ is Brody hyperbolic, and therefore, by the theorems of Brody [25, 3.6.3] and Bloch [25, 3.9.19], it follows that $X_{0}$ is Kobayashi hyperbolic. Let $d_{\bullet}$ denote the Kobayashi pseudodistance. The decreasing property of $d_{\bullet}$ shows that $d_{X}\geq\alpha^{*}(d_{X_{0}})$. On the other hand, $d_{X}$ vanishes on $W$, and so $d_{X}=\alpha^{*}(d_{X_{0}})$, which is a metric on $X_{0}$. This can be seen by lifting chains of discs in $X_{0}$ to $X$ and using the fact that $d_{V}$ is identically zero so that liftings of discs abutting to distinct points in $V$ can be connected inside $V$ by adding discs without increasing the pseudodistance. We leave the easy details to the reader. ∎
@@
If a projective manifold (Calabi–Yau or hyperkähler) $V$ exists which is not hyperbolic but contains no rational curve and no image of an Abelian variety, the preceding construction gives a general type manifold $X$ embedded in $A\times V$ which is not hyperbolic but violates Lang conjecture (the conditional Theorem 1.1).
@@
Choose $V$ to be a projective manifold containing a Zariski dense entire curve in the construction of $X\subset B:=A\times V$ in the proof of Theorem 1.10. The arguments given there apply (except that now $V$ may contain rational curves so that $X$ is only of general type and does not need to have a nef and big cotangent bundle). We can now only assert that $d_{X}\geq\alpha^{*}(d_{X_{0}})$, here not that we have equality. But at least, $d_{X}$ is a distance on $X\setminus W$. ∎
@@
If $d_{V}$ vanishes identically, then we still have, as above, the equality $d_{X}=\alpha^{*}(d_{X_{0}})$. If $V$ is ‘special’ in the sense of [5], it is conjectured that $d_{V}$ vanishes identically. Hence, it is conjectured that this equality holds.
Question 3.4.
@@
Let us consider the product $X\times Z$ where $X\subset A\times V$ is one of the examples constructed in the proof of Theorems 1.10 or 1.11, and $Z$ is canonically polarised and hyperbolic. The exceptional set (in the sense of S. Lang) of $X\times Z$ is then a finite number of copies of $V\times Z$.
4. Remarks on the existence of metrics with quasi-negative holomorphic sectional curvature
Recall that the Kobayashi–Lang conjecture predicts that a compact hyperbolic Kähler manifold is canonically polarised, which is the case if it admits a Kähler metric with negative holomorphic sectional curvature ([41, 39]). This remains true for particular classes of pluriclosed manifolds (see [3, 31]).
It is well known that canonically polarised manifolds need not be hyperbolic: Fermat hypersurfaces in $\mathbf{P}^{n}$ of degree $d>n+1$ support rational curves. Hirzebruch [18] and, more recently, Sarem [38] (building on [20]), have produced canonically polarised manifolds with no rational curves. These examples are toroidal compactifications of ball quotients; hence, they admit elliptic curves and are not hyperbolic.
The examples constructed in Theorem 1.10, however, show that canonically polarised manifolds may have arbitrary simple Abelian variety as their exceptional loci.
The examples constructed by Hirzebruch [18] and Sarem [38] are also of interest since they provide examples of compact Kähler manifolds with (a Kähler metric of) quasi-negative holomorphic sectional curvature [^7] The holomorphic sectional curvature $\text{HSC}:T_{X}^{1,0}\longrightarrow\mathbf{R}$ is said to be quasi-negative if $\text{HSC}(\xi_{x})\leq 0$ for all $\xi_{x}\in T_{X,x}^{1,0}$ and all $x\in X$, and there exists some $x\in X$ such that $\max_{\xi_{x}\in T_{X,x}^{1,0}\setminus\{0\}}\text{HSC}(\xi_{x})<0$., but no (Hermitian) metric of negative holomorphic sectional curvature. Since these examples are compactifications of ball quotients $\mathbf{B}^{n}/\Gamma$, it is immediate that the (holomorphic sectional) curvature is negative on $\mathbf{B}^{n}/\Gamma$. Hence, the difficulty in establishing the quasi-negative curvature lies in showing that the curvature is globally non-positive.
@@ 4. Remarks on the existence of metrics with quasi-negative holomorphic sectional curvature
The projective manifolds $X$ constructed in Theorem 1.10 have Kähler metrics of non-positive bisectional curvature and pseudoconvex Finsler metrics of quasi-negative holomorphic sectional curvature. If $\dim X=2$, then $X$ has no Kähler metric of non-positive sectional curvature.
@@
Maintaining the previous notation, let us recall that $X$ is a submanifold of $A\times V$, where $A$ and $V$ are Abelian varieties. We also constructed $X$ as the blow-up of $X_{0}:=\psi^{-1}(W)\subset A$ at $\psi^{-1}(v)$. We will write $\pi$ for the blow-up map. Let $\delta$ be the metric on $X$ induced from a flat Kähler metric on $A\times V$. From the curvature decreasing property of the holomorphic bisectional curvature, it follows immediately that $X$ has non-positive holomorphic bisectional curvature.
Let $X_{0}^{\circ}$ denote the regular locus of $X_{0}$. We note that $X_{0}$ has ample cotangent bundle, and Kobayashi’s construction [24] yields a pseudoconvex Finsler metric with negative curvature on $X_{0}$. We leave the reader to consult [24] for the definitions, where they are well presented. We note, however, that what we call ‘pseudoconvex’, Kobayashi refers to as ‘convex’ in [24].
Let $U:=\pi^{-1}(X_{0}^{\circ})\subseteq X$. Let $\eta\in\mathcal{C}_{0}^{\infty}(U)$ be a smooth cut-off function, compactly supported in $U$. Let $F$ denote the (pullback to $U$) of the negatively curved pseudoconvex Finsler metric on $X_{0}^{\circ}$. In particular, the holomorphic sectional curvature (interpreted in the sense of Wu [43]) of $\eta F$ is negative on the support of $\eta$. We then apply Wu’s theorem [43] on the holomorphic sectional curvature of the sum of (Finsler) metrics to deduce that the pseudoconvex Finsler metric $\delta+\eta F$ has quasi-negative holomorphic sectional curvature.
The last assertion follows from a result of Zheng [44], which states (in particular) that a canonically polarised surface with a Kähler metric of non-positive sectional curvature is hyperbolic. ∎
From Proposition 4.1 (and motivated by questions in [13, 15]) we naturally propose the following question:
Question 4.2.
@@
The toroidal compactifications of ball quotients considered in [18, 38] have quasi-negative holomorphic bisectional curvature but no Hermitian metrics of negative holomorphic sectional curvature (since they are not hyperbolic!). In this case, the difficulty concentrates on showing that the negatively curved metrics on the ball quotients retain globally non-positive curvature on the compactification (c.f., [20]). The difficulty in our situation is reversed: The examples constructed in Theorem 1.10 are submanifolds of Abelian varieties and, thus, have globally non-positive holomorphic bisectional curvature. Hence, the problem is concentrated on producing strict negativity at some point.

### v2

@@ 2. Statements
Proceed by contradiction and suppose that $f:\mathbf{C}\longrightarrow X$ is an entire curve. Let $V:=\overline{f(\mathbf{C})}^{\text{Zar}}$ denote its Zariski closure. We choose the entire curve such that the dimension $d\geq 1$ of $V$ is minimal. Since we assumed that $X$ does not support any rational curves or images of Abelian varieties, $d>1$. Let $V^{\text{norm}}$ denote the normalization of $V$. Since $V$ is not uniruled, the Kodaira dimension of $V^{\text{norm}}$ is nonnegative, by Conjecture 2.5.
The singularities of $V$ can be resolved to produce a smooth model $V^{\text{smth}}$. Running the minimal model program then yields a terminal good minimal model $V^{\text{term}}$ of $V^{\text{smth}}$. The canonical bundle of the terminal model is nef, and hence, semi-ample by the abundance conjecture. Let $\Phi:V^{\text{term}}\longrightarrow W$ denote the Moishezon–Iitaka map. The generic fiber $V_{w}^{\text{term}}=\Phi^{-1}(w)$ has terminal singularities and vanishing Kodaira dimension. Hence, by the abundance conjecture, the canonical bundle of $V_{w}^{\text{term}}$ is torsion. Since $V$ does not contain any rational curve, $V_{w}^{\text{term}}$ does not contain a Zariski dense set of rational curves. Hence, by Conjecture 2.6, it is a terminal quotient of an Abelian variety. In particular, $V_{w}^{\text{term}}$ contains Zariski dense entire curves. Since the dimension $d$ is chosen to be minimal, we must have $V_{w}^{\text{term}}\simeq V^{\text{term}}$, and therefore, $W$ is a point, and $V^{\text{norm}}$ is a terminal quotient of a simple Abelian variety.
From Lemma 3.1, item 3, we get: $V^{\text{norm}}\simeq V^{\text{term}}$, and so $V^{norm}$ is a terminal undercover of an Abelian variety, contradicting our assumptions. Thus $X$ is (Brody) hyperbolic.
We finally show that $K_{X}$ is ample. Since $X$ does not contain rational curves, $K_{X}$ is nef (by [Mori82]), hence semi-ample (by Conjecture 2.5). If $\Phi:X\to Z$ is the Moishezon-Iitaka fibration, its smooth fibres have torsion canonical bundle and no rational curve, hence are étale undercovers of Abelian varieties. Contradiction unless $X=Z$, which means that $K_{X}$ is ample, as claimed. ∎
@@
In the proof of Theorem 2.7, we only made use of the following weaker form of Conjecture 2.6: Let $V^{\text{norm}}$ be a normal projective variety with no rational curves. Let $V^{\text{term}}$ be a terminal model with torsion canonical bundle, and write $\psi:V^{\text{term}}\dashrightarrow V^{\text{norm}}$ for the rational map. Let $\text{Exc}(\psi)$ and $\Gamma(\psi)$ denote the exceptional locus and graph of $\psi$, respectively. If all rational curves of $V^{\text{term}}$ are contained in the image of $\text{Exc}(\psi)$ under the projection $\Gamma(\psi)\longrightarrow V^{\text{term}}$, then $V^{\text{term}}$ is a terminal quotient of an Abelian variety.
5. Examples: The Proof of Theorems  2.8 and 2.9.
@@
In this section, we prove Theorems 2.8 and 2.9, producing examples of canonically polarised manifolds with exceptional locus (in the sense of Lang) any given simple Abelian variety, or more generally manifolds with Zariski dense entire curves.
@@
Choose for $V$ a simple Abelian variety. The projective manifold $X$ constructed above is canonically polarised, does not contain any rational curve and any elliptic curve if $d\geq 2$. Any entire curve in $X$ is Zariski dense in some of the exceptional divisors $V_{j}\cong V$ of $\alpha:X\longrightarrow X_{0}$ which lie over $\psi^{-1}(v)$ and are the only subvarieties of $X$ not of general type. Moreover, the cotangent bundle of $X$ is nef and big, and the Kobayashi pseudodistance of $X$ is a (non-degenerate) distance on $X\setminus W$, and vanishes on $W$, the union of the $V_{j}^{\prime}s$.
@@
Since $A$ is simple, any of its subvarieties, and so in particular $X_{0}$, are of general type. So is thus the manifold $X$. Since neither $A$ nor $V$ carries rational curves, $K_{X}$ is ample (nef by [Mori82], ample by [Kaw91]). Since $A$ is simple, $X_{0}$ does not carry any rational curve or Abelian subvariety. Thus $X_{0}$ is hyperbolic, and any entire curve of $X$ must be contained in the exceptional locus of $\alpha:X\longrightarrow X_{0}$, which is the union of the $V_{j}^{\prime}s$. Since each $V_{j}$ is isomorphic to the simple Abelian variety $V$, each entire curve of $X$ is contained and Zariski dense in one of the $V_{j}^{\prime}s$ (by Bloch’s theorem).
We prove that the cotangent bundle of $X$ is nef and big. By [Deb05], Proposition 5, it is sufficient for this to check that $X\subset B=A\times V$ is non-degenerate, that is, that the natural difference map $\delta:X\times X\to B$ defined by $\delta(x,y):=x-y$ is generically finite, using the group law on $B$, whatever the origin. Since $A$ and $V$ are simple and not isogeneous (they have different dimensions), the only Abelian subvarieties of $B$ are $A$ and $V$, and so we only need to check that for both projections from $B$ to $A$ and $V$, the image of $X$ is either everything, or of the same dimension as $X$. But this is obvious, since $X$ maps surjectively onto $V$, and birationally to $X_{0}$ in $A$.
We now prove that the Kobayashi pseudodistance on $X$ is a distance on $X\setminus W$. To this end, we observe that $X_{0}\subset A$ is Brody hyperbolic, and therefore, by the theorems of Brody [Kob98, 3.6.3] and Bloch [Kob98, 3.9.19], it follows that $X_{0}$ is Kobayashi hyperbolic. Let $d_{\bullet}$ denote the Kobayashi pseudodistance. The decreasing property of $d_{\bullet}$ shows that $d_{X}\geq\alpha^{*}(d_{X_{0}})$. On the other hand, $d_{X}$ vanishes on $W$, and so $d_{X}=\alpha^{*}(d_{X_{0}})$, which is a metric on $X_{0}$. This can be seen by lifting chains of discs in $X_{0}$ to $X$ and using the fact that $d_{V}$ is identically zero, so that liftings of discs abutting to distinct points in $V$ can be connected inside $V$ by adding discs, without increasing the pseudodistance. We leave the easy details to the reader. ∎
@@
If a projective manifold (Calabi–Yau or hyperkähler) $V$ exists which is not hyperbolic, but contains no rational curve and no image of an Abelian variety, the preceding construction gives a general type manifold $X$ embedded in $A\times V$ which is not hyperbolic, but violates Lang conjecture (the conditional Theorem 2.7).
@@
Choose $V$ to be a projective manifold containing a Zariski dense entire curve in the construction of $X\subset B:=A\times V$ in the proof of Theorem 2.8. The arguments given there apply (except that now $V$ may contain rational curves, so that $X$ is only of general type, and does not need to have a nef and big cotangent bundle). Indeed, $X\to X_{0}\subset A$ is a blow-up in finitely many points and a smooth model of $X_{0}$. Since $A$ is simple, $X_{0}$ and its smooth models are of general type, and do not contain any entire curve (by Bloch’s theorem again). By brody theorem, $d_{X_{0}}$ is a metric, and $d_{X}\geq\alpha^{*}(d_{X_{0}})$. But we do not know whether we have equality. At least, $d_{X}$ is a distance on $X\setminus W$. ∎
@@
If $d_{V}$ vanishes identically, then we still have, as above, the equality $d_{X}=\alpha^{*}(d_{X_{0}})$. If $V$ is ‘special’ in the sense of [Ca04], it is conjectured that $d_{V}$ vanishes identically. Hence, it is conjectured that this equality holds. If $V$ contains a dense (for the metric topology) entire curve, then $d_{V}$ vanishes, and $d_{X}=\alpha^{*}(d_{X_{0}})$.
Question 5.4.
@@
Remark $5.$
Let us consider the product $X\times Z$ where $X\subset A\times V$ is one of the examples constructed in the proof of Theorems 2.8 or 2.9, and $Z$ is canonically polarised and hyperbolic. The exceptional set (in the sense of S. Lang) of $X\times Z$ is then a finite number of copies of $V\times Z$.
6. Remarks on metrics with quasi-negative holomorphic sectional curvature
Recall that the Kobayashi–Lang conjecture predicts that a compact hyperbolic Kähler manifold is canonically polarised, which is the case if it admits a Kähler metric with negative holomorphic sectional curvature ([WY16, TY17]). This remains true for particular classes of pluriclosed manifolds (see [BS24, LeSt21]).
It is well known that canonically polarised manifolds need not be hyperbolic: Fermat hypersurfaces in $\mathbf{P}^{n}$ of degree $d>n+1$ support rational curves. Hirzebruch [Hi84] and, more recently, Sarem [Sar23] (building on [HuSc96]), have produced canonically polarised manifolds with no rational curves. These examples are toroidal compactifications of ball quotients; hence, they admit elliptic curves and are not hyperbolic.
@@ 6. Remarks on metrics with quasi-negative holomorphic sectional curvature
The referee showed us easier examples of non-hyperbolic surfaces with no rational curves, but with ample canonical bundle: the second symmetric power of a bi-elliptic, non-hyperelliptic curve $C$ of genus at least $3$. The exceptional locus is then a finite union of elliptic curves.
The examples constructed in Theorem 2.8 however show that canonically polarised manifolds may have arbitrary simple Abelian variety as their exceptional loci.
The examples constructed by Hirzebruch [Hi84] and Sarem [Sar23] are also of interest since they provide examples of compact Kähler manifolds with (a Kähler metric of) quasi-negative holomorphic sectional curvature [^7] The holomorphic sectional curvature $\text{HSC}:T_{X}^{1,0}\longrightarrow\mathbf{R}$ is said to be quasi-negative if $\text{HSC}(\xi_{x})\leq 0$ for all $\xi_{x}\in T_{X,x}^{1,0}$ and all $x\in X$, and there exists some $x\in X$ such that $\max_{\xi_{x}\in T_{X,x}^{1,0}\setminus\{0\}}\text{HSC}(\xi_{x})<0$., but no (Hermitian) metric of negative holomorphic sectional curvature. Since these examples are compactifications of ball quotients $\mathbf{B}^{n}/\Gamma$, it is immediate that the (holomorphic sectional) curvature is negative on $\mathbf{B}^{n}/\Gamma$. Hence, the difficulty in establishing the quasi-negative curvature lies in showing that the curvature is globally non-positive.
@@
The projective manifolds $X$ constructed in Theorem 2.8 have Kähler metrics of nonpositive bisectional curvature and pseudoconvex Finsler metrics of quasi-negative holomorphic sectional curvature. If $\dim X=2$, then $X$ has no Kähler metric of nonpositive sectional curvature.
@@
Maintaining the previous notation, let us recall that $X$ is a submanifold of $A\times V$, where $A$ and $V$ are Abelian varieties. We also constructed $X$ as the blow-up of $X_{0}:=\psi^{-1}(W)\subset A$ at $\psi^{-1}(v)$. We will write $\pi$ for the blow-up map. Let $\delta$ be the metric on $X$ induced from a flat Kähler metric on $A\times V$. From the curvature decreasing property of the holomorphic bisectional curvature, it follows immediately that $X$ has nonpositive holomorphic bisectional curvature.
Let $X_{0}^{\circ}$ denote the regular locus of $X_{0}$. We note that $X_{0}$ has ample cotangent bundle, and Kobayashi’s construction [Kob75] yields a pseudoconvex Finsler metric with negative curvature on $X_{0}$. We leave the reader to consult [Kob75] for the definitions, where they are well presented. We note, however, that what we call ‘pseudoconvex’, Kobayashi refers to as ‘convex’ in [Kob75].
Let $U:=\pi^{-1}(X_{0}^{\circ})\subseteq X$. Let $\eta\in\mathcal{C}_{0}^{\infty}(U)$ be a smooth cut-off function, compactly supported in $U$. Let $F$ denote the (pullback to $U$) of the negatively curved pseudoconvex Finsler metric on $X_{0}^{\circ}$. In particular, the holomorphic sectional curvature (interpreted in the sense of Wu [Wu73]) of $\eta F$ is negative on the support of $\eta$. We then apply Wu’s theorem [Wu73] on the holomorphic sectional curvature of the sum of (Finsler) metrics to deduce that the pseudoconvex Finsler metric $\delta+\eta F$ has quasi-negative holomorphic sectional curvature.
The last assertion follows from a result of Zheng [Z95] which states (in particular) that a canonically polarised surface with a Kähler metric of nonpositive sectional curvature is hyperbolic. ∎
From Proposition 6.2 (and motivated by questions in [Di, DiTr19]) we naturally propose the following question:
Question 6.3.
@@
The toroidal compactifications of ball quotients considered in [Hi84, Sar23] have quasi-negative holomorphic bisectional curvature, but no Hermitian metrics of negative holomorphic sectional curvature (since they are not hyperbolic!). In this case, the difficulty concentrates in showing that the negatively curved metrics on the ball quotients retain globally nonpositive curvature on the compactification (c.f., [HuSc96]). The difficulty in our situation is reversed: The examples constructed in Theorem 2.8 are submanifolds of Abelian varieties and thus, have globally nonpositive holomorphic bisectional curvature. Hence, the problem is concentrated on producing strict negativity at some point.